# What Do Privacy Advertisements Communicate to Consumers?

### v1

@@ 3.2. Recruitment and Screening
We recruited participants for our study using convenience sampling, snowball sampling and Prolific, a research-oriented online crowdworking platform (Prolific, 2023). To reduce privacy and security priming, we presented the study as a study on “Consumer Perspective on Tech Advertising.” We omitted any words related to privacy or security in the study title and purpose, letting the campaigns deliver relevant privacy components naturally. We also did not mention any company name related to the campaigns when we advertised the study. Participants were compensated \$1.25 USD for completing a screening survey and those who subsequently participated in the 25-minute interview were compensated \$10 USD for their time.
We designed four treatments to test the five campaigns: Apple, Samsung, Twitter, and WhatsApp-DuckDuckGo. With the exception of Twitter which was an online game, all other campaigns were short videos ranging from 30 to 95 seconds. WhatsApp and DuckDuckGo ads did not introduce actionable privacy features that we could ask participants about and were the shortest of all video campaigns, with their combined video length being roughly the same as that of Apple and Samsung. In the interest of maintaining similar interview length across all treatments, WhatsApp and DuckDuckGo were grouped together.
We required eligible participants to be aged 18 or above, fluent in English, and located within the United States (the primary target location for the versions of the campaigns we tested). With the exception of Twitter, whose game was primarily intended for its users, all remaining video campaigns were ads likely aimed at the general public, both users and non-users. Thus we developed a screening survey, aimed at discerning participants’ product or service usage that we used to assign participants to treatment groups. We used purposive sampling to select from among those who qualified for our study and assign balanced treatment groups. The screening survey is in Appendix A.1.
For Apple, Samsung and WhatsApp-DuckDuckGo treatment groups, we screened for 50% users and 50% non-users of the relevant product or service. The Twitter treatment group was comprised of 100% Twitter users with some form of online and/or digital gaming experience. We defined users as those who utilize the product or service featured in the campaign at least once every week. Due to low interview response rate from eligible DuckDuckGo participants, we relaxed the qualifying user condition from using DuckDuckGo at least once a week to at least once a month. We also asked participants about their technical backgrounds and distributed those with more technical backgrounds evenly across treatment groups.
@@
Initially, we screened 181 participants from Prolific using our screening survey, and invited 133 participants who were eligible. Only 24 participants responded to the invite, including seven who rejected the invite, and three who missed their appointment. Thus, only 14 participants completed both the survey and the interview. Due to low response rate for interview invitations with Prolific participants, we then began recruiting via snowball sampling using email invitations. The first author emailed acquaintances unaware with the research topic using the same information as had been presented to Prolific participants. These acquaintances were also invited to refer others who might be interested. Using this method, we recruited 10 participants. In total, 24 participants were recruited for this study, split evenly into four treatment groups. We ensured that for each treatment group, there were no more than three participants who indicated they had technical expertise. All participants who were recruited through Prolific were paid via the platform, whereas snowball participants were compensated at the same rate, but via Amazon gift cards.
3.3. Interview Procedure
@@
We developed a semi-structured interview procedure, with an estimated completion time of 25 minutes. All interviews were conducted over the Zoom video conferencing platform and recorded and transcribed by Zoom. Participants were asked to keep their cameras off throughout the interview. Interviews were conducted between March 31 to April 19, 2023.
@@
First, we asked about participants’ general impression of the companies, then screen-shared the advertisements into Zoom for participants to watch. For Twitter DataDash, we sent each participant a Twitter login details and the game link. Once participants had the game open, they were asked to share their screen over Zoom and given 8 minutes of DataDash playtime while the interviewer observed them. In the other treatments, participants spent up to 95 seconds viewing a video advertisement. We asked participants a series of questions aimed at gauging their awareness of privacy features that were shown in the campaign, as well as discussing any privacy and security concepts mentioned by participants. We asked participants for their impression of the companies again after exposure to the campaigns to see whether the campaign had an impact on this. Then, we asked participants to describe how actionable they thought the privacy advice provided was, and self-rate their motivation and confidence in being able to complete the steps shown in the ad or game. Next, we moved to the actionability exercise where we asked all participants except those in the WhatsApp-DuckDuckGo treatment to find a specific setting that was shown in the ad or game, with a total activity time of 3 minutes. Finally, for ad treatment groups, we re-exposed participants to the ad and asked for concluding remarks and final observations.
@@
Since the WhatsApp and DuckDuckGo videos did not contain other actionable steps aside from switching to the advertised service, we did not include the actionability section of the interview for that treatment, but instead repeated the entire interview procedure again for the second ad. The order of the advertisements shown (WhatsApp first, then DuckDuckGo, or vice versa) was randomized.
@@
The main purpose of the actionability section of the interview, was to assess perceived ease of actionability, measured through participants’ description of what steps they thought the advertisement had recommended, as well as self-ranked motivation and confidence. We shared an iPhone 12 Pro, a Galaxy A13 smartphone, or the Twitter home page over Zoom. Participants were asked to direct the interviewer on where to go to find the appropriate setting. Prior to the study, we had restored both smartphones’ desktops to factory default desktop settings.
@@
Video advertising campaigns are generally designed with the assumption that viewers will see ads multiple times (Landa, 2016). Therefore, in the three ad treatments we re-exposed participants to the ads again at the end of the interview to see if this would generate any new insights.
@@
Before recruiting participants for our study, we conducted eight pilot interviews, two per treatment group. We made a number of revisions to our protocol based on issues discovered during the pilots. Pilot interviews are not included in our results and analysis.
@@ 3.4. Data Analysis
Interview transcripts were generated automatically using Zoom’s built-in transcription function and manually reviewed with the voice recordings by a member of the research team. Two researchers performed qualitative analysis for each participant interview transcript using Template Analysis, a style of thematic analysis that combines both inductive and deductive coding (King, 2024; Brooks and King, 2014). Template Analysis emphasizes hierarchical coding without specific prescription regarding the number of levels required and what levels represent. To analyze the data, both researchers first familiarized themselves with the data and the first researcher created an initial codebook (the template) and coded all the data. Then, the second researcher reviewed and refined the codes applied by the first researcher. Both researchers met over multiple sessions to discuss any disagreements in the coding and adjusted the codebook (added, updated, and deleted codes) until they reached a satisfiable version of the codebook, agreeable to both researchers. Our coding started with broader themes, such as “attitudes,” “awareness,” and “actionability,” which encompass narrower themes, such as “organization impression pre/post-exposure,” to more specific ones, such as “positive,” “negative,” etc. For the actionability section (task), after the research team agreed on a standard criteria (also provided in Appendix A.6) to evaluate task completion, one researcher recorded whether participants completed the activity, then recorded the participants’ self-rated scores regarding motivation and confidence. A diagram summarizing the coding hierarchy is provided in Appendix A.5. The codebook along with the specific interview question(s) from which each set of codes were distilled is provided in Appendix A.6.
@@ 3.5. Ethical Considerations
This study protocol was approved by [Anonymous Institution]’s Institutional Review Board (IRB). Before completing the screening survey, we asked all participants to read and acknowledge their acceptance of the consent form. We asked all interview participants to keep their cameras off. Before participants in the Twitter treatment were asked to share their screen over Zoom we reminded them to avoid accidental information sharing. Since the Twitter game actionability activity required users to login and navigate Twitter’s account settings, participants were provided with an account by the researcher and were not asked to login using their personal accounts. A member of the research team removed participant names from transcripts along with any personal information participants shared in the interviews.
@@ 4. Results
In this section, we discuss the main finding of this study. We use the first letter from each company name to indicate which treatment group the participants are in, followed by the participant number. While our study is qualitative, we provide numbers of occurrence of qualitative themes to highlight the most or least common themes.
@@ 4.1. Participant Demographics
We interviewed 24 participants, including 13 males, and 11 females. Participants’ ages ranged from 18 to 65 years old. Of the 24 participants, 10 indicated they have a technical background. All Apple and Samsung participants who were asked about their smartphone use said they actively use smartphone apps and browse the Internet using their smartphone. All but one DuckDuckGo participants said they actively browse the web and perform search queries using search engines. For more details about participant demographics, refer to Appendix Section A.4.
@@ 4.2. Attitudes and Impressions of Companies
We asked participants about their impressions of the companies relevant to their treatment group both before and after viewing the videos or playing the game (Q.5 & Q.14 in Appendix A.2). We reviewed the transcripts of these responses and categorized the sentiments they expressed before watching the ads or playing the game into five categories: positive, neutral, negative, mixed-positive (more positive than negative), and mixed-negative (more negative than positive). We classified participants who mentioned an equal number of positive and negative points based on whether the first point they mentioned was positive or negative. We used the same five codes to analyze responses after viewing the ads or playing the game, with two additional codes: positive-plus and negative-plus, to categorize responses that indicated increased degree of positivity or negativity. We then compared each participant’s sentiment before and after viewing the ad or playing the game, and classified the change as: more positive, more negative, or remained the same.
Our results suggest that most of the campaigns we studied were associated with brands that our participants already viewed fairly positively, with the exception of Twitter. Exposure to information about privacy features further improved some participants’ impressions about these companies, although for a few participants, the campaigns worsened their impressions as they viewed claims about privacy in the campaigns with skepticism. The Twitter game also worsened the impression of some participants due to both poor game play experience and unclear messaging.
@@ 4.2.1. Attitude and Impression of Companies Before Exposure
Top reasons for having a positive attitude towards an organization across all groups were related to having a good product or having a privacy-focused product. Some participants who were not users of a product still remarked that they thought the product was good (A2, S2, S6, WD6 for WhatsApp and DuckDuckGo). S2, a Samsung non-user, remarked: “I think that they come out with top notch products. I think I would be almost as happy having a Samsung Android type phone as I would having an iPhone.” DuckDuckGo received the most mentions of privacy-focused aspects as a reason for positive attitudes towards the company, mentioned by both users and non-users. For example, WD5 said they have “positive feelings about [DuckDuckGo]” due to “the privacy features.”
Twitter had the most negative impressions, with two negatives (T1, T2) and one mixed-negative impression (T4). Participants mentioned false news (T1, T4) and lack of moderation on the platform (T1, T2). From T2’s perspective, “what Twitter used to be, was not perfect, but tolerable. And now it’s just a single rich person’s plaything.” It is interesting to note that when discussing their views of Twitter, three participants (T2, T4, T5) either directly mentioned Twitter owner Elon Musk, or implied that the leadership change had some impact on their impression of the organization. One Apple participant (A4) mentioned high prices as the reason for their negative impression of the company. A4 explained, “An iMac will cost on average, about twice as much … as a given windows machine [of] approximately equivalent capability.”
@@ 4.2.2. Attitude and Impression of Companies After Exposure
There were also five participants (from Twitter, DuckDuckGo, and WhatsApp) who, after being exposed to the campaign, had a more negative view of the organization. Skepticism was the top reason for having a worsened impression, mentioned by three participants (WD2, WD4 for DuckDuckGo, WD4 for WhatsApp). Some of those participants questioned the truthfulness of the message and whether the organizations will deliver on the messages the ads portrayed. WD2 wanted to do more research on DuckDuckGo, because “well, how does it make money?” In addition, prior knowledge of negative news about the company such as data breaches could also lead to skepticism, with WD4 being skeptical about both WhatsApp and DuckDuckGo due to his exposure to news about those companies. WD4 felt “Slightly more negative” about DuckDuckGo due to “what I’ve read online because they’re still reinforcing this, you know, one this silver bullet of search engines,” likely referring to the Microsoft DuckDuckGo tracker incident (Fadilpašić, 2022). WD4 admitted that he“wasn’t diligent enough about following up what happened with WhatsApp about a year ago,” likely referring to the WhatsApp data breach (Espósito, 2021).
For the two Twitter participants who had a more negative view of the organization after being exposed to the campaign, their reasoning stemmed from a poor experience with the DataDash game itself. T1 remarked that “privacy and safety issues are a very important issue, and and it seems like this, this game didn’t really do a good job of taking it like very seriously … I think there are a lot more better ways to educate than this game … I think if they did it again have like more real life implications, and it was more tied to real examples of what can happen with the privacy and data.” T5 had a much more sinister take on what the game represented, concluding that Twitter “probably [has] no issue selling our data and like exploiting our privacy.” From T5’s understanding, “the bones [represented] the data pieces, or like information … the dog was probably just someone scooping up data pieces.”
@@ 4.3. Communicating Privacy
To assess how well each campaign communicated privacy, we asked participants to describe the campaign they had just viewed or experienced, how they felt towards the ad or game, and what they perceived as the purpose of the campaigns (Q.7, Q.8, and Q.9). From these responses, we explored how well the campaigns increased awareness of privacy features and issues, how well participants understood the privacy metaphors used in the campaign, and their understanding of the purpose of the campaign.
Overall, our findings suggest that participants were able to grasp broader privacy themes communicated in the campaigns. In general, after being exposed to the campaign, participants noticed at least one main privacy feature showcased, and their perception of the privacy metaphors mostly aligned with that of the research team. The WhatsApp campaign, which appeared to be most successful at communicating a clear message was short, communicated a single feature (end-to-end encryption) and used a simple metaphor (pigeon to represent unencrypted SMS). On the other hand, Twitter’s DataDash game was the only campaign that faced considerable challenges in communicating its privacy feature, with some participants failing to recognize the privacy connection until we explicitly asked about it.

### v2

@@ 3.2. Recruitment and Screening
We recruited participants for our study using convenience snowball sampling and Prolific, a research-oriented online crowdworking platform (Prolific, 2023). To reduce privacy and security priming, we presented the study as a study on “Consumer Perspective on Tech Advertising.” We omitted any words related to privacy or security in the study title and purpose, letting the campaigns deliver relevant privacy components naturally. We also did not mention any company name related to the campaigns when we advertised the study. Participants were compensated 1.25 USD for completing a screening survey and those who subsequently participated in the 25-minute interview were compensated 10 USD for their time.
We designed four treatments to test the five campaigns: Apple, Samsung, Twitter, and WhatsApp-DuckDuckGo. Except for Twitter which was an online game, all other campaigns were short videos ranging from 30 to 95 seconds. WhatsApp and DuckDuckGo ads did not introduce actionable privacy features that we could ask participants about, and were the shortest of all video campaigns, with their combined video length being roughly the same as that of the Apple and Samsung videos. In the interest of maintaining similar interview lengths across all treatments, WhatsApp and DuckDuckGo were grouped together.
We required eligible participants to be aged 18 or above, fluent in English, and located within the United States (the primary target location for the versions of the campaigns we tested). We administered a screening survey that collected demographic information and information about participants’ product or service usage. We used purposive sampling to select from among those who qualified for our study and assign balanced treatment groups. The screening survey is in Appendix A.1.
For Apple, Samsung, and WhatsApp-DuckDuckGo treatment groups, we screened for 50% users and 50% non-users of the relevant product or service since those campaigns appeared to be aimed at both users and non-users of the advertised products and services. The Twitter treatment group was comprised of 100% Twitter users with some form of online and/or digital gaming experience. We defined users as those who utilize the product or service featured in the campaign at least once every week. Due to the low interview response rate from eligible DuckDuckGo participants, we relaxed the qualifying user condition from using DuckDuckGo “At least once a week” to “At least once a month.” We also asked participants about their technical backgrounds and distributed those with technical backgrounds evenly across treatment groups. We defined technical background as a degree or work experience in one or more of the following areas: Computer Science, Information Systems, Information Technology, and Computer Engineering.
@@
Initially, we screened 181 participants from Prolific using our screening survey and invited 133 eligible participants. Only 24 participants responded to the invitation, including seven who rejected the invitation, and three who missed their appointment. Thus, only 14 participants completed both the survey and the interview. Due to the low response rate from Prolific participants, we also recruited via snowball sampling using email invitations. The first author emailed acquaintances unaware of the research topic using the same information as had been presented to Prolific participants. These acquaintances were also invited to refer others who might be interested. Using this method, we recruited 10 participants. In total, we recruited 24 participants for this study, split evenly into four treatment groups. We ensured that for each treatment group, there were no more than three participants who indicated they had technical expertise. All participants who were recruited through Prolific were paid via the platform, whereas snowball participants were compensated at the same rate, but via Amazon gift cards.
$3.$ Interview Procedure
@@
We developed a semi-structured interview procedure, with an estimated completion time of 25 minutes. All interviews were conducted over the Zoom video conferencing platform and recorded and transcribed by Zoom. Participants were asked to keep their cameras off throughout the interview. Interviews were conducted between March 31 and April 19, 2023.
@@
First, we asked about participants’ general impression of the companies, then screen-shared the advertisements into Zoom for participants to watch. For Twitter DataDash, we sent each participant Twitter login details and the game link. Once participants had the game open, they were asked to share their screen over Zoom and given 8 minutes of DataDash playtime while the interviewer observed them. In the other treatments, participants spent up to 95 seconds viewing a video advertisement. We asked participants a series of questions aimed at gauging their awareness of privacy features that were shown in the campaign, as well as discussing any privacy and security concepts mentioned by participants. We asked participants for their impression of the companies again after exposure to the campaigns to see whether the campaign had an impact on this. Then, we asked participants to describe how actionable they thought the privacy advice provided was, and self-rate their motivation and confidence in being able to complete the steps shown in the ad or game. Next, we moved to the actionability exercise where we asked all participants except those in the WhatsApp-DuckDuckGo treatment to find a specific setting that was shown in the ad or game, with a total activity time of 3 minutes. Finally, for ad treatment groups, we re-exposed participants to the ad and asked for concluding remarks and final observations.
@@
Since the WhatsApp and DuckDuckGo videos did not contain other actionable steps aside from switching to the advertised service, we did not include the actionability section of the interview for that treatment, but instead repeated the entire interview procedure for the second ad. The order of the advertisements shown (WhatsApp first, then DuckDuckGo, or vice versa) was randomized.
@@
The main purpose of the actionability section of the interview was to assess perceived ease of actionability, measured through participants’ description of what steps they thought the advertisement had recommended, as well as self-ranked motivation and confidence. We shared an iPhone 12 Pro, a Galaxy A13 smartphone, or the Twitter home page over Zoom. Participants were asked to direct the interviewer on where to go to find the appropriate settings. Before the study, we had restored both smartphones’ desktops to factory default desktop settings.
@@
Video advertising campaigns are generally designed with the assumption that viewers will see ads multiple times (Landa, 2016). Therefore, in the three ad treatments, we re-exposed participants to the ads again at the end of the interview to see if this would generate any new insights.
@@
Before recruiting participants for our study, we conducted eight pilot interviews, two per treatment group. We made several revisions to our protocol based on issues discovered during the pilots. Pilot interviews are not included in our results and analysis.
@@ 3.4. Data Analysis
Interview transcripts were generated automatically using Zoom’s built-in transcription function and manually reviewed with the voice recordings by a member of the research team. Two researchers performed qualitative analysis for each participant interview transcript using Template Analysis, a style of thematic analysis that combines both inductive and deductive coding (King, 2024; Brooks and King, 2014). Template Analysis emphasizes hierarchical coding without specific requirements regarding the number of levels and what levels represent. To analyze the data, both researchers first familiarized themselves with the data. Next, the first researcher created an initial codebook (the template) and coded all the data. Then, the second researcher reviewed and refined the codes applied by the first researcher. Both researchers met over multiple sessions to discuss any disagreements in the coding and adjusted the codebook (added, updated, and deleted codes) until they reached a satisfactory version of the codebook, agreeable to both researchers. Our coding started with broader themes, such as “attitudes,” “awareness,” and “actionability,” which encompass narrower themes, such as “organization impression pre/post-exposure,” to more specific ones, such as “positive,” “negative,” etc. For the actionability section (task), after the research team agreed on standard criteria (also provided in Appendix A.6) to evaluate task completion, one researcher recorded whether participants completed the activity, and then recorded the participants’ self-rated scores regarding motivation and confidence. A diagram summarizing the coding hierarchy is provided in Appendix A.5. The codebook along with the specific interview question(s) from which each set of codes were distilled is provided in Appendix A.6.
@@ 3.5. Ethical Considerations
This study protocol was approved by CMU’s Institutional Review Board (IRB). Before completing the screening survey, we asked all participants to read and acknowledge their acceptance of the consent form. We asked all interview participants to keep their cameras off. Before participants in the Twitter treatment were asked to share their screen over Zoom we reminded them to avoid accidental information sharing. Since the Twitter game actionability activity required users to log in and navigate Twitter’s account settings, participants were provided with an account by the researcher and were not asked to log in using their personal accounts. A member of the research team removed participant names from transcripts along with any personal information participants shared in the interviews.
@@ 4. Results
In this section, we discuss the main findings of this study. We use the first letter from each company name to indicate which treatment group the participants were in, followed by the participant number. While our study is qualitative, we provide frequencies of themes to illustrate how common each theme was in our dataset.
@@ 4.1. Participant Demographics
We interviewed 24 participants, including 13 males, and 11 females. Participants’ ages ranged from 18 to 65 years old. Of the 24 participants, 10 indicated they have a technical background. All Apple and Samsung participants who were asked about their smartphone use said they actively use smartphone apps and browse the Internet using their smartphones. All but one DuckDuckGo participant said they actively browse the web and perform search queries using search engines. For more details about participant demographics, refer to Table 2 in Appendix A.4.
@@ 4.2. Attitudes and Impressions of Companies
We asked participants about their impressions of the companies relevant to their treatment group both before and after viewing the videos or playing the game (Q.5 & Q.14 in Appendix A.2). We reviewed the transcripts of these responses and categorized the sentiments they expressed before watching the ads or playing the game into five categories: positive, neutral, negative, mixed-positive (more positive than negative), and mixed-negative (more negative than positive). We classified participants who mentioned an equal number of positive and negative points based on whether the first point they mentioned was positive or negative. We used the same five codes to analyze responses after viewing the ads or playing the game, with two additional codes: positive-plus and negative-plus, to categorize responses that indicated an increased degree of positivity or negativity. We then compared each participant’s sentiment before and after viewing the ad or playing the game, and classified the change as: more positive, more negative, or remained the same.
Our results suggest that most of the campaigns we studied were associated with brands that our participants already viewed fairly positively, except Twitter. Exposure to information about privacy features further improved some participants’ impressions of these companies, although, for a few participants, the campaigns worsened their impressions as they viewed claims about privacy in the campaigns with skepticism. The Twitter game also worsened the impression of some participants due to both poor gameplay experience and unclear messaging.
@@ 4.2.1. Attitude and Impression of Companies Before Exposure
The top reasons for having a positive attitude towards an organization across all groups were related to having a good product or having a privacy-focused product. Some participants who were not users of a product still remarked that they thought the product was good (A2, S2, S6, WD6 for WhatsApp and DuckDuckGo). S2, a Samsung non-user, remarked: “I think that they come out with top-notch products. I think I would be almost as happy having a Samsung Android type phone as I would having an iPhone.” DuckDuckGo received the most mentions of privacy-focused aspects as a reason for positive attitudes towards the company, mentioned by both users and non-users. For example, WD5 said they have “positive feelings about [DuckDuckGo]” due to “the privacy features.”
Twitter had the most negative impressions, with two negatives (T1, T2) and one mixed-negative impression (T4). Participants mentioned false news (T1, T4) and lack of moderation on the platform (T1, T2). From T2’s perspective, “what Twitter used to be, was not perfect, but tolerable. And now it’s just a single rich person’s plaything.” It is interesting to note that when discussing their views of Twitter, three participants (T2, T4, T5) either directly mentioned Twitter owner Elon Musk, or implied that the leadership change had some impact on their impression of the organization. One Apple participant (A4) mentioned high prices as the reason for their negative impression of the company. A4 explained: “an iMac will cost on average, about twice as much … as a given Windows machine [of] approximately equivalent capability.”
@@ 4.2.2. Attitude and Impression of Companies After Exposure
There were also five participants (from Twitter, DuckDuckGo, and WhatsApp) who, after being exposed to the campaign, had a more negative view of the organization. Skepticism was the top reason for having a worsened impression, mentioned by three participants (WD2, WD4 for DuckDuckGo, WD4 for WhatsApp). Some of those participants questioned the truthfulness of the message and whether the organizations would deliver the messages the ads portrayed. WD2 wanted to do more research on DuckDuckGo, because “well, how does it make money?” In addition, prior knowledge of negative news about the company such as data breaches could also lead to skepticism. For example, WD4 was skeptical about both WhatsApp and DuckDuckGo due to his exposure to news about those companies. WD4 felt “Slightly more negative” about DuckDuckGo due to “what I’ve read online because they’re still reinforcing this … this silver bullet of search engines,” likely referring to the Microsoft DuckDuckGo tracker incident (Fadilpas̆ić, 2022). WD4 admitted that he “wasn’t diligent enough about following up what happened with WhatsApp about a year ago,” likely referring to the WhatsApp data breach (Espósito, 2021).
For the two Twitter participants who had a more negative view of the organization after being exposed to the campaign, their reasoning stemmed from a poor experience with the DataDash game itself. T1 remarked that “privacy and safety issues are a very important issue, and and it seems like this, this game didn’t really do a good job of taking it like very seriously … I think there are a lot more better ways to educate than this game … I think if they did it again have like more real-life implications, and it was more tied to real examples of what can happen with the privacy and data.” T5 had a much more sinister take on what the game represented, concluding that Twitter “probably [has] no issue selling our data and like exploiting our privacy.” From T5’s understanding, “the bones [represented] the data pieces, or like information … the dog was probably just someone scooping up data pieces.”
@@ 4.3. Communicating Privacy
To assess how well each campaign communicated about privacy, we asked participants to describe the campaign they had just viewed or experienced, how they felt toward the ad or game, and what they perceived as the purpose of the campaigns (Q.7, Q.8, Q.9). From these responses, we explored how well the campaigns increased awareness of privacy features and issues, how well participants understood the privacy metaphors used in the campaigns, and their understanding of the purpose of the campaign.
Overall, our findings suggest that participants were able to grasp broader privacy themes communicated in the campaigns. In general, after being exposed to the campaign, participants noticed at least one main privacy feature showcased, and their perception of the privacy metaphors mostly aligned with that of the research team. The WhatsApp campaign, which appeared to be most successful at communicating a clear message, was short, communicated a single feature (end-to-end encryption), and used a simple metaphor (pigeon to represent unencrypted SMS). On the other hand, Twitter’s DataDash game was the only campaign that faced considerable challenges in communicating its privacy feature, with some participants failing to recognize the privacy connection until we explicitly asked about it.